OpenStudy (anonymous):

Please help , what's 2x^-1/3+5=15

OpenStudy (anonymous):

Step 1: Multiply both sides by 3x.

OpenStudy (anonymous):

15x+2=45x 15x+2−45x=45x−45x(Subtract 45x from both sides) −30x+2=0 −30x+2−2=0−2(Subtract 2 from both sides) −30x=−2 -30x/-30 = -2/-30(Divide both sides by -30) x=1/5

OpenStudy (anonymous):

Um... you're welcome